OpenStudy (anonymous):

true or false? the graph of the equation y=13 is a horizontal line?

OpenStudy (mathlegend):

@apeach01 did you try graphing it?

OpenStudy (mathlegend):

If that is saying you are on the y-axis... and you go up to 13... will the line be horizontal?

OpenStudy (anonymous):

there is no graph or anything other than what I wrote...true or false...the graph of the equation y=13 is a horizontal line? that is all it says and has...

OpenStudy (anonymous):

True. It is a straight line that goes horizontally through the y-axis.

OpenStudy (anonymous):

for every value of x, y=13 

OpenStudy (mathlegend):

Good so you know exactly what you are talking about @ape

OpenStudy (anonymous):

oh my gosh I feel so stupid! that is really all the question was asking? I thought there was more too it! thank you

OpenStudy (mathlegend):

@apeach01 the graph above is how it would look

OpenStudy (anonymous):

:) sometimes we just expect them to be throwing us a fast curve ball high and outside when it is just a slow pitch right in the center of the strike zone.

OpenStudy (anonymous):

Find the intercepts for the graph of the equation...-3x+3y=3 I am trying to help a friend and cannot remember how to work out this problem!! i know the formulas rise over run and y2-y1 over x2-x1 and I also know y=mx-b but I cannot seem to figure this one out!! I need to be refreshed